# Cobblestone: A Divide-and-Conquer Approach for Automating Formal Verification

### v3

@@ Abstract.
copyright: none
@@ 1. Introduction
Bugs in software systems can be costly and dangerous. In 2022, poor software quality cost the US economy \$2.41 trillion (Krasner22, ), and bugs can bring down critical, global systems (Menn24, ). Formal verification using proof assistants, such as Coq (coq, ) or Lean (Moura21, ), is a promising method of improving software quality. Formal verification can mathematically prove the absence of entire classes of bugs, providing strong guarantees for the correctness of critical software systems. And formal verification is highly effective: A study (Yang11a, ) of C compilers found bugs in every tested compiler, including LLVM (Lattner04, ) and GCC (Stallman12, ), but not in the formally verified (in Coq) portions of CompCert (Leroy09, ).
@@
But, formal verification requires specifying desired properties, writing mathematical proofs of the properties, and machine checking those proofs using the proof assistant. Writing these proofs requires significant expertise and manual effort. For example, the proofs verifying CompCert are 8 times longer than the functional code (Leroy06, ), and even small changes to the software can require heavy proof editing (Ringer21, ). While hundreds of large software systems have been verified (Ringer21, ), like the sel4 microkernel (Klein09, ; Murray13, ) and the CakeML compiler (kumar2014cakeml, ), and formal verification has seen industrial success, e.g., at Airbus France (Souyris14, ), Google, and Mozilla (Erbsen19, ; Jacobs20, ), most software today is not verified due to the high manual cost.
@@
Recent work has aimed to reduce the cost of formal verification by using machine learning to synthesize proofs (Sanchez20, ; First20oopsla, ; First22icse, ; Sanchez-Stern23toplas, ; Paliwal20, ; Yang19, ; First23fse, ). However, these approaches, even when combined with an SMT-solver-based hammer (Czajka18, ) can only prove one third of the desired properties on a large benchmark of open-source Coq projects (First22icse, ). The advent of large language models (LLMs) has improved performance in this space (lu2024proofautomationlargelanguage, ; thompson2025rango, ; thakur2024an, ). In this paper, we aim to occupy the happy medium between two common approaches to proof synthesis.
@@
Some work uses LLMs to generate whole proofs at once, without using the intermediate proof state produced by partial proofs (First23fse, ; zhang2023getting, ). For instance, PALM (lu2024proofautomationlargelanguage, ) samples an LLM for a whole proof attempt and then uses repair mechanisms. The success of whole-proof-based approaches relies on the powerful reasoning abilities of proprietary LLMs, such as OpenAI and Anthropic models, as they must construct proofs that follow viable high-level plans. However, these approaches fail to provide the LLM with access to the intermediate state, which contains important context to aid in reasoning.
@@
Other approaches perform tactic-by-tactic search, building a proof one step at a time and allowing the LLM to condition the next step generation based on access to the intermediate state (Yang19, ; First20oopsla, ; Sanchez20, ; Yang23, ; Blaauwbroek24ICML, ; Jiang2022Thor, ). However, navigating the complex search space of possible proofs is challenging since the local state steers the search without consideration for a high-level proof plan.
@@
These approaches represent two extremes on the spectrum of proof synthesis granularity. Our work is a middle ground between these two extremes, incorporating the advantages of whole-proof generation with leveraging intermediate states where this generation fails. To achieve this middle ground, we make a fundamental insight: we can leverage the tree structure of whole-proof attempts to localize valid and invalid subproofs. We can keep valid subproofs and recurse on points of failure where we regain access to the state. This insight motivated the design of our tool Cobblestone, a novel LLM approach to proof synthesis.
@@
We present Cobblestone, which uses a divide and conquer algorithm to combine an LLM’s high-level planning ability with the advantages of local exploration. Cobblestone starts by tackling the entire goal, automatically identifying which parts of its proof are working and which are not, progressively reducing the granularity of its attack from the whole proof to individual proof steps by recursing, if necessary.
@@
Cobblestone uses an LLM to generate whole proofs, analyzes the proofs using its novel fail-safe execution to decompose the proof into subgoals, and determines which subgoals’ subproofs work and which fail. It then attempts to repair the failing subproofs, and, if necessary, recurses on the remaining, unproven subgoals to build a whole, correct proof. Because Cobblestone divides the goal into smaller subgoals, it is more effective at using automated theorem proving tools such as hammers; while a hammer may fail to prove the overall goal, it is more likely to succeed on smaller, decomposed goals. Unlike prior work POETRY (wang2024proving, ), Cobblestone does not require fine-tuning an LLM to identify proof substructure and locations to recurse, making it easily extensible to use different LLMs (requiring only minor changes to API calls).
We implement Cobblestone for the Coq proof assistant, and evaluate it on GPT-4 and two other LLMs. We evaluate on subsets of four benchmarks: CoqGym (Yang19, ), a set of open-source Coq projects from GitHub used for evaluating prior proof synthesis tools (Sanchez20, ; First20oopsla, ; First22icse, ; Sanchez-Stern23toplas, ; Yang19, ); and coq-wigderson (10.1145/3618305.3623600, ), coq-bb5 (bb5github, ), and PnVRocqLib (pnvgithub, ), three recent projects deliberately chosen to be outside of the time range of GPT-4’s training cutoff to control for the impact of pretraining data leakage on LLMs.
We compare Cobblestone to prior non-LLM-based proof synthesis tools, LLM-based tools, and LLM baselines of our own creation. Prior RNN-based proof synthesis tool, Proverbot9001 (Sanchez20, ) proves 17% of the CoqGym subset and 10% of the coq-wigderson subset, while SMT-solver-based CoqHammer (Czajka18, ) proves 30% and 27%, respectively. We implement a baseline LLM-based approach that uses chain-of-thought reasoning and show that it proves 25% and 19%, respectively. Meanwhile, Cobblestone, automatically proves 48% of the CoqGym subset and 38% of the coq-wigderson subset. When prior LLM-based tool PALM (lu2024proofautomationlargelanguage, ) is run multiple times to match the token usage of our approach, Cobblestone outperforms it on the coq-wigderson, coq-bb5, and PnVRocqLib subsets, and is highly complementary with it on the CoqGym subset. Therefore, Cobblestone complements prior work by proving different theorems they do not, so that together, they prove significantly more.
@@
  • A method for executing a proof to localize proof errors and separate working and failing parts of a proof.
@@
To ensure reproducibility of our results and enable others to build on our work, we will make all code, experimental scripts, and data publicly available (anonCobblestoneReplicationPackage, ).
@@ 2. The Cobblestone Approach
To formally verify a software system, in addition to writing the code for the system, an engineer needs to write a mathematical theorem formalizing the desired property, and a high-level proof (proof script) that the property holds. Cobblestone automatically writes the proof script. Given code and a theorem stating a property (e.g., in_adj_exists in Figure 2), Cobblestone tries to synthesize a proof using a divide and conquer approach, as shown in Figure 1. First, Cobblestone tries synthesizing a proof using CoqHammer (Section 2.3). If that fails, Cobblestone samples an LLM for candidate solutions (Section 2.4). If none of the candidates prove the theorem entirely, Cobblestone uses a novel fault localization method to isolate errors while noting which subgoals are successfully proven (Section 2.5). Next, Cobblestone keeps successful parts of candidate solutions and recurses to find proofs for the yet unproven subgoals (Section 2.6), ultimately combining them into a proof of the original theorem.
@@ 2.1. Illustrative Example
.111from graph.v in the coq-wigderson project Figure 2 lists a definition (lines 1–4), the in_adj_exists theorem statement (lines 6–8), and a human-written proof for the theorem (lines 9–15). This theorem states that for all graphs g and all nodes i, j, if i is adjacent to j in g, then g contains v, a set of nodes adjacent to j, and that i $\in$ v.
@@
The rest of the proof script is split into two subproofs, one for each subgoal, each starting with a bullet (-). The first bullet on line 13 invokes CoqHammer using the hammer tactic. CoqHammer (Czajka18, ) attempts to automatically prove a goal using SMT solvers and proof reconstruction procedures. If hammer succeeds, the goal is proven. Otherwise, the tactic produces an error.
@@ 2.2. Cobblestone on the Illustrative Example
Instead of naïvely trying to prove in_adj_exists again, as prior work has tried (First23fse, ), Cobblestone recurses on each subgoal, attempting to prove it instead ([Uncaptioned image] and [Uncaptioned image]). In [Uncaptioned image], Cobblestone tries using a hammer, which fails, and then prompts an LLM for a proof of the subgoal “M.find j g = Some (adj g j)”. This results in a proof with an error, which Cobblestone temporarily fixes by adding the admit tactic. The proof with admit, a proof in progress, is passed as an argument to another recursive call. Here, Cobblestone changes the proof in progress to a working proof by replacing admit with the hammer tactic. In [Uncaptioned image], Cobblestone is able to prove “S.In i (adj g j)” using only hammer.
@@ 2.3. Applying CoqHammer
Cobblestone combines neural-based proof synthesis and SMT-based proof synthesis by invoking CoqHammer (Czajka18, ). In Figure 1, the hammer icon ([Uncaptioned image]) shows how Cobblestone uses CoqHammer in two ways. First, at the start of Cobblestone’s execution (and in subsequent recursive calls), Cobblestone invokes CoqHammer on the goal, which sometimes produces a working proof. Second, in recursive calls, Cobblestone may be passed a proof in progress, which consists of working tactics followed by an “admit”. For example, “unfold adj, M.find in *. admit.” in Figure 4 [Uncaptioned image] is a proof in progress. Cobblestone replaces admit with hammer, and checks if this proves the subgoal. Attempting to repair proofs in progress like this allows Cobblestone to use CoqHammer on a variety of goals, increasing the frequency of success.
@@ 2.4. Sampling LLM for Whole Proofs
Central to Cobblestone’s proof synthesis is using an instruction-tuned LLM (ouyang2022training, ) to sample whole proofs. Cobblestone uses a prompt that consists of two strings — a setup prompt with high-level directions about the task the LLM should perform, and a sampling prompt with details specific to the current theorem. In LLM terminology, the setup prompt is called a “system message” and the sampling prompt is called a “user message,” though Cobblestone generates these prompts automatically, without a user.
The setup prompt directs the LLM to produce a whole proof for the provided theorem. The sampling prompt contains 5 types of information: (1) the theorem statement to prove, (2) the current proof state, (3) the definitions for all identifiers mentioned in the theorem statement and proof state that are not in the standard library, (4) contextual information, and (5) optional reasoning, which is used as part of chain-of-thought prompting. The setup prompt is formatted as below, with each section preceded by a section header. Cobblestone constructs the sampling prompt fully automatically using the coq-serapy library (Sanchez20, ) to interact with Coq.
@@
Chain-of-thought (Wei2022-qv, ), a popular prompting framework that prompts an LLM to provide natural language reasoning before generating its final response, has been shown to help improve LLM performance. Cobblestone implements chain-of-thought by first prompting the LLM with a modified setup prompt that instructs it to generate reasoning, and then prompting it a second time, with the generated reasoning added to the [REASONING] section of the sampling prompt. When Cobblestone samples an LLM without chain-of-thought prompting, this extra section is omitted.
Prior work has shown that the use of diverse data in models can increase the proving power of a proof-synthesis approach (First22icse, ). Accordingly, to increase variation in Cobblestone’s generated proofs, Cobblestone samples the LLM with 4 variations of the same prompt in 2 different dimensions — with and without contextual information and with and without chain-of-thought reasoning. Cobblestone benefits from this approach because the Coq theorem prover can serve as an oracle of proof correctness to pick a correct proof from among multiple samples. Our replication package (anonCobblestoneReplicationPackage, ) includes all prompts that were used as part of our evaluation.
@@ 2.5. Localizing Errors with Fail-Safe Mode
Algorithm 1 describes fail-safe mode, which takes a proof script broken into prefixes and subproofs as input, modifies it to allow execution after errors, and annotates subproofs, subsubproofs, etc. as “correct” or “in progress.” First, it executes each tactic in the prefix. If a tactic results in an error, that tactic and everything after it is replaced with admit, and the prefix is annotated as “in progress.” Note that inserting admits into proofs in progress is temporary; Cobblestone will remove these in a recursive call.
@@
The first complication is that the proof can encounter an error while executing a prefix tactic after it decomposes the original goal. Concretely, destruct a might successfully decompose the goal into subgoals, but then, apply nonexistent can fail. In this case, instead of just replacing the apply nonexistent with an admit, fail-safe mode searches for a shorter prefix that leads to multiple subgoals, which we call a decomposing prefix. For example, if the sequence intros. destruct a. leads to multiple subgoals, then fail-safe mode keeps it as the prefix and creates a bullet with an admit placeholder for each subgoal, as shown on the right.
The second complication is that a decomposing tactic can produce more or fewer subgoals than the number of bullets that follow. On the right, there are two bullets that follow destruct a. If destruct a produces fewer subgoals than two, fail-safe mode trims the number of bullets; if destruct a produces more than two, fail-safe mode adds the appropriate number of bullets with admit placeholders. This guarantees that the number of subproofs that appear syntactically in the proof matches the number of subgoals generated by the prefix.

### v4

@@ Abstract.
journalyear: 2026copyright: ccconference: 2026 IEEE/ACM 48th International Conference on Software Engineering; April 12–18, 2026; Rio de Janeiro, Brazilbooktitle: 2026 IEEE/ACM 48th International Conference on Software Engineering (ICSE ’26), April 12–18, 2026, Rio de Janeiro, Brazildoi: 10.1145/3744916.3773178isbn: 979-8-4007-2025-3/2026/04
[…]
by
@@ 1. Introduction
Bugs in software systems can be costly and dangerous. In 2022, poor software quality cost the US economy \$2.41 trillion (Krasner, 2022), and bugs can bring down critical, global systems (Menn and Gregg, 2024). Formal verification using proof assistants, such as Coq (The Coq Development Team, 2017) or Lean (de Moura and Ullrich, 2021), is a promising method of improving software quality. Formal verification can mathematically prove the absence of entire classes of bugs, providing strong guarantees for the correctness of critical software systems. And formal verification is highly effective: A study (Yang et al., 2011) of C compilers found bugs in every tested compiler, including LLVM (Lattner and Adve, 2004) and GCC (Stallman, 2012), but not in the formally verified (in Coq) portions of CompCert (Leroy, 2009).
@@
But, formal verification requires specifying desired properties, writing mathematical proofs of the properties, and machine checking those proofs using the proof assistant. Writing these proofs requires significant expertise and manual effort. For example, the proofs verifying CompCert are 8 times longer than the functional code (Leroy, 2006), and even small changes to the software can require heavy proof editing (Ringer, 2021). While hundreds of large software systems have been verified (Ringer, 2021), like the sel4 microkernel (Klein et al., 2009; Murray et al., 2013) and the CakeML compiler (Kumar et al., 2014), and formal verification has seen industrial success, e.g., at Airbus France (Souyris, 2014), Google, and Mozilla (Erbsen et al., 2019; Jacobs and Beurdouche, 2020), most software today is not verified due to the high manual cost.
@@
Recent work has aimed to reduce the cost of formal verification by using machine learning to synthesize proofs (Sanchez-Stern et al., 2020; First et al., 2020; First and Brun, 2022; Sanchez-Stern et al., 2023; Paliwal et al., 2020; Yang and Deng, 2019; First et al., 2023). However, these approaches, even when combined with an SMT-solver-based hammer (Czajka and Kaliszyk, 2018) can only prove one third of the desired properties on a large benchmark of open-source Coq projects (First and Brun, 2022). The advent of large language models (LLMs) has improved performance in this space (Lu et al., 2024; Thompson et al., 2025; Thakur et al., 2024a).
@@
In this paper, we aim to occupy the happy medium between two common approaches to proof synthesis. Some work uses LLMs to generate whole proofs at once, without using the intermediate proof state produced by partial proofs (First et al., 2023; Zhang et al., 2023). For instance, PALM (Lu et al., 2024) samples an LLM for a whole proof attempt and then uses repair mechanisms. The success of whole-proof-based approaches relies on the powerful reasoning abilities of proprietary LLMs, such as OpenAI and Anthropic models, as they must construct proofs that follow viable high-level plans. However, these approaches fail to provide the LLM with access to the intermediate state, which contains important context to aid in reasoning.
@@
Other approaches perform tactic-by-tactic search, building a proof one step at a time and allowing the LLM to condition the next step generation based on access to the intermediate state (Yang and Deng, 2019; First et al., 2020; Sanchez-Stern et al., 2020; Yang et al., 2023; Blaauwbroek et al., 2024; Jiang et al., 2022). However, navigating the complex search space of possible proofs is challenging since the local state steers the search without consideration for a high-level proof plan.
@@
These approaches represent two extremes on the spectrum of proof synthesis granularity. Our work is a middle ground between these two extremes, incorporating the advantages of whole-proof generation with leveraging intermediate states where this generation fails. To achieve this middle ground, we make a fundamental insight: we can leverage the tree structure of whole-proof attempts to localize valid and invalid subproofs. We can keep valid subproofs and recurse on points of failure where we regain access to the state. This insight motivated the design of our tool Cobblestone, a novel LLM-based approach to proof synthesis.
[…]
@@
We present Cobblestone, which uses a divide and conquer algorithm to combine an LLM’s high-level planning ability with the advantages of local exploration. Cobblestone starts by tackling the entire goal, automatically identifying working and broken parts of the proof, reducing the granularity of its attack from the whole proof to individual proof steps by recursing, if necessary.
[…]
@@
Cobblestone uses an LLM to generate whole proofs, analyzes the proofs using its novel fail-safe execution to decompose the proof into subgoals, and determines which subgoals’ subproofs work and which fail. It then attempts to repair the failing subproofs, and, if necessary, recurses on the remaining, unproven subgoals to build a whole, correct proof. Because Cobblestone divides the goal into smaller subgoals, it is more effective at using automated theorem proving tools such as hammers; while a hammer may fail to prove the overall goal, it is more likely to succeed on smaller, decomposed goals. Unlike prior work POETRY (Wang et al., 2024a), Cobblestone does not require fine-tuning an LLM to identify proof substructure and locations to recurse, making it easily extensible to use different LLMs (requiring only minor changes to API calls).
We implement Cobblestone for the Coq proof assistant, and evaluate it on GPT-4 and two other LLMs. We evaluate on subsets of four benchmarks: CoqGym (Yang and Deng, 2019), a set of open-source Coq projects from GitHub used for evaluating prior proof synthesis tools (Sanchez-Stern et al., 2020; First et al., 2020; First and Brun, 2022; Sanchez-Stern et al., 2023; Yang and Deng, 2019); and coq-wigderson (Phipathananunth, 2023), coq-bb5 (mxdys, 2025), and PnVRocqLib (PnVDiscord, [n. d.]), three recent projects deliberately chosen to be after GPT-4’s training cutoff to control for the impact of pretraining data leakage on LLMs.
We compare Cobblestone to prior non-LLM-based proof synthesis tools, LLM-based tools, and LLM baselines of our own creation. Prior RNN-based proof synthesis tool, Proverbot9001 (Sanchez-Stern et al., 2020) proves 17% of the CoqGym subset and 10% of the coq-wigderson subset, while SMT-solver-based CoqHammer (Czajka and Kaliszyk, 2018) proves 30% and 27%, respectively. We implement a baseline LLM-based approach that uses chain-of-thought reasoning and show that it proves 25% and 19%, respectively. Meanwhile, Cobblestone, automatically proves 48% of the CoqGym subset and 38% of the coq-wigderson subset. When prior LLM-based tool PALM (Lu et al., 2024) is run multiple times to match the token usage of our approach, Cobblestone outperforms it on the coq-wigderson, coq-bb5, and PnVRocqLib subsets, and is highly complementary with it on the CoqGym subset. Therefore, Cobblestone complements prior work by proving different theorems they do not, so that together, they prove significantly more.
@@
  • Fail-safe mode, a method for executing a proof to localize proof errors and separate working and failing parts of a proof.
@@
To ensure reproducibility of our results and enable others to build on our work, we make all code, experimental scripts, and data publicly available (Kasibatla et al., 2025).
@@ 2. The Cobblestone Approach
To formally verify software, in addition to writing the code for a system, an engineer needs to write a mathematical theorem formalizing the desired property, and a high-level proof (proof script) that the property holds. Cobblestone automatically writes the proof script. Given code and a theorem (e.g., in_adj_exists in Figure 2), Cobblestone synthesizes a proof using a divide and conquer approach, as shown in Figure 1. First, Cobblestone tries synthesizing a proof using CoqHammer (Section 2.3). If that fails, Cobblestone samples an LLM for candidate solutions (Section 2.4). If none of the candidates prove the theorem entirely, Cobblestone uses a novel fault localization method to isolate errors while noting which subgoals are successfully proven (Section 2.5). Next, Cobblestone keeps successful parts of candidate solutions and recurses to find proofs for the yet unproven subgoals (Section 2.6), ultimately combining them into a proof of the original theorem.
@@ 2.1. Illustrative Example
.111from graph.v in the coq-wigderson project Figure 2 lists a definition (lines 1–4), the theorem statement (lines 6–8), and a human-written proof (lines 9–15). This theorem states that for all graphs g and all nodes i, j, if i is adjacent to j in g, then g’s structure must contain v.
@@
The rest of the proof script is split into two subproofs, one for each subgoal, each starting with a bullet (-). The first bullet on line 13 invokes CoqHammer using the hammer tactic. CoqHammer (Czajka and Kaliszyk, 2018) attempts to automatically prove a goal using SMT solvers and proof reconstruction procedures. If hammer succeeds, the goal is proven. Otherwise, the tactic produces an error.
@@ 2.2. Cobblestone on the Illustrative Example
Instead of naïvely trying to prove in_adj_exists again, as prior work has tried (First et al., 2023), Cobblestone recurses on each subgoal, attempting to prove it instead ([Uncaptioned image] and [Uncaptioned image]). In [Uncaptioned image], Cobblestone tries using a hammer, which fails, and then prompts an LLM for a proof of the subgoal “M.find j g = Some (adj g j)”. This results in a proof with an error, which Cobblestone temporarily fixes by adding the admit tactic. The proof with admit, a proof in progress, is passed as an argument to another recursive call. Here, Cobblestone changes the proof in progress to a working proof by replacing admit with the hammer tactic. In [Uncaptioned image], Cobblestone is able to prove “S.In i (adj g j)” using only hammer.
@@ 2.3. Applying CoqHammer
Cobblestone combines neural-based proof synthesis and SMT-based proof synthesis by invoking CoqHammer (Czajka and Kaliszyk, 2018). In Figure 1, the hammer icon ([Uncaptioned image]) shows how Cobblestone uses CoqHammer in two ways. First, at the start of Cobblestone’s execution, Cobblestone invokes CoqHammer on the goal, which sometimes produces a working proof. Second, in recursive calls, Cobblestone may be passed a proof in progress, which consists of working tactics followed by an “admit”. For example, “unfold adj, M.find in *. admit.” in Figure 4 [Uncaptioned image] is a proof in progress. Cobblestone replaces admit with hammer, and checks if this proves the subgoal. Attempting to repair proofs in progress like this allows Cobblestone to use CoqHammer on a variety of goals, increasing the frequency of success.
@@ 2.4. Sampling LLM for Whole Proofs
Central to Cobblestone’s proof synthesis is using an instruction-tuned LLM (Ouyang et al., 2022) to sample whole proofs. Cobblestone’s prompt consists of two parts—a system message with high-level directions about the task the LLM should perform, and a user message333“user message” is terminology used by the OpenAI API. Cobblestone generates these prompts automatically, without a user. with details specific to the current theorem.
The system message directs the LLM to produce a whole proof for the provided theorem. The user message contains 5 types of information: (1) the theorem statement to prove, (2) the current proof state, (3) the definitions for all identifiers mentioned in the theorem statement and proof state that are not in the standard library, (4) contextual information, and (5) optional reasoning, which is used as part of chain-of-thought prompting. The user message is formatted as below, with each section preceded by a section header. Cobblestone constructs the user message fully automatically using the coq-serapy library (Sanchez-Stern et al., 2020) to interact with Coq.
@@
Chain-of-thought (Wei et al., 2022a), a popular prompting technique, prompts an LLM to provide natural language reasoning before generating its final response, and has been shown to help improve LLM performance. Cobblestone implements chain-of-thought by first prompting the LLM with a modified system message that instructs it to generate reasoning, and then prompting it a second time, with the generated reasoning added to the [REASONING] section of the user message. When Cobblestone samples an LLM without chain-of-thought prompting, this extra section is omitted.
Prior work has shown that the use of diverse data in models can increase the proving power of a proof-synthesis approach (First and Brun, 2022). Accordingly, to increase variation in Cobblestone’s generated proofs, Cobblestone samples the LLM with 4 variations of the same prompt in 2 different dimensions—with and without contextual information and with and without chain-of-thought reasoning. Cobblestone benefits from this approach because the Coq theorem prover can serve as an oracle of proof correctness to pick a correct proof from among multiple samples. Our replication package (Kasibatla et al., 2025) includes all prompts that were used as part of our evaluation.
@@ 2.5. Localizing Errors with Fail-Safe Mode
Algorithm 1 describes fail-safe mode, which takes a proof script broken into prefixes and subproofs as input, modifies it to allow execution after errors, and annotates subproofs, subsubproofs, etc. as “correct” or “in progress.” First, it executes each tactic in the prefix. If a tactic results in an error, that tactic and everything after it is replaced with admit [^4] Note that inserting admits into proofs in progress is temporary; Cobblestone will remove these in a recursive call., and the prefix is annotated as “in progress.”
@@
The first complication is that the proof can encounter an error while executing a prefix tactic after it decomposes the original goal. Concretely, “destruct a” might successfully decompose the goal into subgoals, but then, “apply nonexistent” can fail. In this case, instead of just replacing the “apply nonexistent” with an admit, fail-safe mode searches for a shorter prefix that leads to multiple subgoals, which we call a decomposing prefix. For example, if the sequence “intros. destruct a.” leads to multiple subgoals, then fail-safe mode keeps it as the prefix and creates a bullet with an admit placeholder for each subgoal, as shown on the right.
The second complication is that a decomposing tactic can produce more or fewer subgoals than the number of bullets that follow. On the right, there are two bullets that follow “destruct a”. If “destruct a” produces fewer subgoals than two, fail-safe mode trims the number of bullets; if “destruct a” produces more than two, fail-safe mode adds the appropriate number of bullets with admit placeholders. This guarantees that the number of subproofs that appear syntactically in the proof matches the number of subgoals generated by the prefix.